given the following reaction, N2 (g) + 3H2 (g) -> 2NH3 (g), how many liters of H2 are required to produce 47 moles of NH3
1
Expert's answer
2018-08-15T04:12:04-0400
N2 (g) + 3 H2 (g) -> 2 NH3 (g)
V (H2) / 3*22.4 L = 47 moles / 2 moles
molar volume Vm = 22.4 L
V (H2) =3*22.4 L* 47 moles / 2 moles = 1579.2 L